Four point charges, each with Q = 4.5 uC are arranged at the corners of a...

  1. Four point charges, each with Q = 4.5 uC are arranged at the corners of a square of edge length 1.5 m. What is the electric potential at the center of the square?
Answer #1

V=kQ/r

So

V=4kQ/(L/sqrt(2))=4sqrt(2)kQ/L=152735.065Volts
Answer #2

Potential of all charges are added as scalar quantities
therefore potential from one charge V = \sqrt{2 }kQ/L

so from 4 charges = 4\sqrt{2 }kQ/L= 4*1.414* 9*10^9*4.5*10^-6/1.5 =152. 712 *10^3 V=1.527*10^5V